in order to have $10,000 in 5 years, how much would a person have to invest if the money will earn 12% interest compounded semiannually?
Answer by mananth(16946)

Principal P = x
Amount= 10000
years=n 5
compounded 2 times a year
Rate = 12.00 6.000
Amount = P*((n+r)/n)^n

10000 = x *( 1 + r/n )^ n


10000 = P ( 1 + 0.060 / 2 ) ^ 2 * 5

10000 = P( 1.06000 )^ 10

10000 /( 1.06 )^ 10 = P

P= 5583.95